OpenStudy (anonymous):

Find all solutions for Sin (2θ) + 3 cos θ = 0

OpenStudy (anonymous):

start with \[2\sin(x)\cos(x)+3\cos(x)=0\] and then it should be easier

OpenStudy (anonymous):

i.e. "double angle" formula \[\cos(x)(2\sin(x)+3)=0\] etc

OpenStudy (anonymous):

satelite first step was very good, we need only one function in order to solve for x, \[2\sin thetacos \theta+3\cos \theta=0\]

OpenStudy (anonymous):

sustract both sides by 3costheta

OpenStudy (anonymous):

\[2\sin \theta \cos \theta= -3\cos\]

OpenStudy (anonymous):

divide now both sides by costheta

OpenStudy (anonymous):

\[2\sin \theta=3\]

OpenStudy (anonymous):

@spcsean can u find ur way from here?